OpenStudy (anonymous):

Determine whether the infinite geometric series is convergent or divergent. If it is convergent, find its sum. (If the series is divergent, enter DIVERGENT.) 1- 1/6 + 1/36 - 1/216 + . . .

OpenStudy (anonymous):

\[\begin{align*}1-\frac{1}{6}+\frac{1}{36}-\frac{1}{216}+\cdots&=\left(-\frac{1}{6}\right)^0+\left(-\frac{1}{6}\right)^1+\left(-\frac{1}{6}\right)^2+\left(-\frac{1}{6}\right)^6+\cdots\\ &=\sum_{n=0}^\infty \left(-\frac{1}{6}\right)^n \end{align*}\]

OpenStudy (anonymous):

How can you tell whether a geometric series converges?

OpenStudy (anonymous):

So what would the sum be?

OpenStudy (perl):

we use a formula to find the sum of a geometric series

OpenStudy (perl):

sum = a / (1-r), where a = first term. r = ratio . here a = 1 , r = -1/6

OpenStudy (perl):

1/ ( 1 - (-1/6) = 1 / ( 1 + 1/6) = 1/ ( 7/6) = 6/7